OpenStudy (matimaticas):

f^-1(3) when f(x) = 2x + 3, all over 5.

OpenStudy (anonymous):

\[\frac{ \left( 2\left( 3 \right)+3 \right)^-1 }{ 5}\]

OpenStudy (matimaticas):

is that 8 over 5?

OpenStudy (anonymous):

2*3=??

OpenStudy (matimaticas):

6

OpenStudy (anonymous):

6+3=9 (9)^-1/5